Savannah wants $12,000 saved in 6 years to make a down payment on a house. How much money should she invest now at 4.3% compounded quarterly in order to meet her goal?

Answer by Theo(13342)

f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.

in your problem.
f = 12,000
p = what you want to find
r = 4.3% per year / 4 = 1.075% per quarter / 100 = .01075 per quarter.
n = 6 years * 4 = 24 quarters.

formula becomes 12,000 = p * (1+.01075) ^ 24
solve for p to get:
p = 12,000 / (1+.01075) ^ 24 = 9283.916931.

confirm by replacing p with that in the original equation to get:
f = 9283.916931 ^ 24 = 12,000

this confirms the value for p is correct.